OpenStudy (anonymous):

Simplify: mn^-4/p^0q^-2 a.) mq^2/p^0n^4 b.) mp/n^4 c.) mn^-4q^2p^0 d.)q^2/mn^4 I know it looks confusing, lol. :o

OpenStudy (anonymous):

@itsmylife can you help mee.(;

OpenStudy (raden):

hint : n^-4 = 1/n^4 1/q^(-2) = q^2 p^0 = 1

OpenStudy (anonymous):

b.?

OpenStudy (raden):

nopes.. it should be : mn^-4/p^0q^-2 = mq^2/p^0n^4 = mq^2/n^4